# CoBRA: A Universal Strategyproof Confirmation Protocol for Quorum-based Proof-of-Stake Blockchains

### v2

@@ 5.1 Lemmas for proving safety
To complete the withdrawal, a second transaction $R_{p,s}$ must be committed to $L_{i}$ at a time $>t^{\prime}+2\Delta^{*}$. Since $L_{i}$ is the first ledger to include $R_{p}$, the same condition applies to any other conflicting ledger, namely the withdrawal can only be finalized after $>t^{\prime}+2\Delta^{*}$. However, since correct validators broadcast valid certificates, every correct validator will witness all conflicting ledgers at some time $t^{*}\leq t+\Delta^{*}\leq t^{\prime}+2\Delta^{*}$. As a result, they will detect equivocation and blacklist $p$, preventing the finalization of $R_{p,s}$. Thus, $p$’s withdrawal cannot succeed in any of the conflicting ledgers.
Strong(-est) Chain Finalization Rule
We provide only sketch proofs for the Strong and Strongest finalization rules (see Section 4.5), and defer the full analysis to Appendix B.
We leverage additional signatures to constrain the number of possible forks and, consequently, the total amount of coins that can be spent in conflicting blocks. Specifically, we bound the coins validators can finalize within $\Delta^{*}$ based on the number of signatures in their local chain. In the most optimistic scenario, with $65n/$ 6 signatures, we show that, regardless of other forks, only one strongest chain can exist (Lemma 10), to which all validators will transition within $\Delta^{*}$. In both cases, we prove that rational validators attempting to equivocate across the other forks can gain at most $D$ coins from conflicting blocks (Lemma 9, Lemma 11). Given our slashing mechanism (Lemma 3), rational validators cannot increase their profit through equivocation, ensuring that the strongest chain finalization rule preserves safety.
@@
Finally, in Lemmas 45, we conclude that rational validators have no incentive to engage in forking attacks, as any potential gain does not exceed their staked amount, which is subject to slashing.
@@
Let us fix any rational validator $p$ and compute its utility to join coalitions forking the system. In Lemmas 2911, we showed that by performing the attack, $p$ can double-spend at most $D$ coins when validators employ i) finalization rule presented in 4, ii) strong ledger finalization rule, iii) strongest finalization rule respectively. According to Lemma 3, $p_{j}$ cannot withdraw its stake $D$ before the conflicting ledgers are detected. Moreover, upon detecting equivocation the system halts and thus misbehaving validators cannot utilize further bribes. Therefore, for any strategy $s_{p}\in S^{p}_{\mathsf{f}}$ and $s_{-p}\in S^{-p}_{\mathsf{f}}$, the utility of $p$ is $u_{p}(s_{p},s_{-p})\leq 0$, while following the protocol yields non-negative utility.
To show that following the protocol weakly dominates the strategy of forking the system, we will construct a configuration where $p$ chooses a strategy $s_{p}\in S^{p}_{\mathsf{f}}$ and fix the rest strategies $s_{-p}\in S^{-p}_{\mathsf{f}}$ such that $u_{p}(s_{p},s_{-p})<0$ (this is sufficient since following the protocol has non-negative utility). To this end, suppose that $p$ participates in a coalition $\mathcal{F}$ consisting of $f$ Byzantine and $\lceil\frac{f+1}{2}\rceil$ rational validators. Moreover, consider a subset $H_{1}$ of $\lceil\frac{f+1}{2}\rceil$ correct validators, and a subset $H_{2}$ distinct to $H_{1}$ that consists of $\lfloor\frac{f+1}{2}\rfloor$ correct validators. Validators in $\mathcal{F}$ construct to conflicting ledgers $T_{1}$ and $T_{2}$ as described in Lemma 8. The utility of $p$ in that case is $u_{p}(s_{p},s_{-p})\leq\frac{D}{\lceil\frac{f+1}{2}\rceil}-D<0$.
@@
For the sake of contradiction assume that two distinct correct validators $p_{1}$, $p_{2}$ finalize the conflicting blocks $b_{1}$ and $b_{2}$ respectively. Validator $p_{1}$ finalized $b_{1}$ after witnessing the respective valid certificate $QC(b_{1})$, and $p_{2}$ finalized $b_{2}$ after witnessing $QC(b_{2})$. By quorum intersection there exists a validator $p$ that is not Byzantine and has voted in both $QC(b_{1})$ and $QC(b_{2})$. Since correct validators do not equivocate, $p$ must be rational. However, according to Lemma 4, for rational validators, following the protocol weakly dominates participating in a coalition that forks the system. So, $p$ will not sign for conflicting blocks and thus we reached a contradiction.
5.2 Lemmas for proving liveness
Since $\Pi$ is $(n,f)$-resilient and $f+1+k\geq n-f$, liveness is ensured as long as rational validators participate in liveness-preserving actions during some intervals. In Lemma 6, we show that there exist windows during which it is a dominant strategy for rational validators to follow the protocol specifications. Building on this, Lemma 7 concludes that some correct validator will successfully propose a block that gathers a valid certificate and finality votes.
@@ 5.2 Lemmas for proving liveness
Forking. From Lemma 4, any strategy in the strategy profile $S^{p}_{f}$ (i.e., where $p$ actively contributes to a fork), is weakly dominated by following the protocol. Therefore, no rational validator will engage in forking attacks.
@@
Suppose the adversary is able to bribe at least one rational validator in every window $W_{m}$ following round $r$. In such windows, at most $2f$ validators can successfully propose valid blocks. Since participation rewards are distributed only when at least $12f{+}$ 1 distinct validators succeed, no rewards are paid out in these bribed windows. Because the adversary’s budget is finite, it is eventually depleted and the adversary cannot bribe any rational validator anymore. Therefore, there exists a future window $W_{k}$ (starting at some round $r^{\prime}\geq r$) such that, for any validator $p$, any strategy $s_{p}\in S_{p}$, and any $s_{-p}\in S_{-p}$ under which $p$ abstains from liveness actions during $W_{k}$, we have $u_{p}(s_{p},s_{-p})\leq 0$.
To show that following the protocol during $W_{k}$ is a weakly dominant strategy, fix any rational validator $p$, and consider two strategies: $s_{1}$, where $p$ follows the protocol during $W_{k}$, and $s_{2}$, where $p$ abstains from liveness actions during $W_{k}$ (we have already established the result for forking in 4). Let $s_{-p}$ be a strategy profile in which at least $2f$ other validators follow the protocol during $W_{k}$. In the combined strategy profile $s_{a}=(s_{1},s_{-p})$, at least $12f+$ 1 validators follow the protocol, satisfying the reward condition, and $p$ receives the corresponding participation rewards. In contrast, under $s_{b}=(s_{2},s_{-p})$, only $2f$ validators follow the protocol, so no rewards are distributed. Hence, $u^{W_{k}}_{p}(s_{a})>u^{W_{k}}_{p}(s_{b})$.
@@
  2. 2.
    $12f+$ 1 finality votes for $b$ or a subsequent block are included in blocks with valid certificates.
@@
According to Lemma 6, for any round $r$, there exists a round $r^{\prime}\geq r$ such that a time window $W_{k}$ begins, during which it is a weakly dominant strategy for rational validators to follow the protocol. Consequently, some block $b$ proposed by an honest validator during $W_{k}$ will obtain a valid quorum certificate $QC(b)$ and receive at least $12f+$ 1 finality votes (either for $b$ or for one of its descendants).
5.3 Final Theorems
Theorem 5.1.
Validators running the composition of CoBRA and $\Pi$ agree in a total order.
@@ 5.3 Final Theorems
Safety. In Lemma 5, we show that validators will never finalize conflicting blocks.
Liveness. Consider any transaction $tx$ witnessed by all correct validators at some round $r$. According to Lemma 7, there is a round $r^{*}\geq r$ where a correct validator $p$ will successfully propose a valid block $b$. If $tx$ has not already been included in a valid block proposed before round $r^{*}$, then, $p$ will include $tx$ at block $b$.
Theorem 5.2.
The composition of CoBRA and $\Pi$ satisfies certifiability.
@@
To this end, consider the inclusion proof $\pi_{tx}$ for the transaction $tx$. Client $c$ has accepted $\pi_{tx}$ only after receiving a proof showing the inclusion of $tx$ in a block $b$, where $b$ has $12f+$ 1 finality votes. At least one correct validator $p$ has committed a finality vote for $b$. Validator $p$ commits finality votes only for blocks included in its transaction ledger.
Client-liveness. For every transaction $tx$ committed to a block $b$ accepted by all correct validators, any correct validator $p_{i}$ can construct a proof $\pi$ as follows. Validator $p_{i}$ sends an inclusion proof $\pi_{tx}$ of $tx$ in $b$ to a client $c$, along with a proof $\pi_{f}(b)$ consisting of $12f+$ 1 finality votes for $b$ or subsequent block, which is feasible according to Lemma 7. Any client $c$ will accept the proof $\pi$ by definition of the client confirmation rule.
Theorem 5.3.
The composition of CoBRA and $\Pi$ is an $(n,k,f^{*})$-resilient SMR protocol.
@@
Validators running the composition of CoBRA and $\Pi$ agree in a total order according to Theorem 5.1. Moreover, in Theorem 5.2 we show that the composition of CoBRA and $\Pi$ satisfies certifiability.
6 Case-Study
These high participation rates are not unique to Cosmos. On Ethereum, the average daily participation has also consistently remained above 99% [18]. Thus, we conclude that CoBRA could be seamlessly integrated into production systems like Cosmos or Ethereum, enhancing system resilience without impacting block confirmation times.
7 Limitations and Extensions
Dynamic Validator Setting
We outline how our approach extends to dynamic settings where new validators may join the system:
  • New validator: A validator that wishes to join the system in round $r$ broadcasts a join request and waits until round $r+2\Delta^{*}$ before fully participating. This waiting period allows the validator to detect any past consistency violations.
  • Existing validator: Fix any existing validator $p$ that receives the join request at some round $r^{\prime}\in\{r,r+\Delta^{*}\}$. It then broadcasts any consistency violations it has observed up to round $r^{\prime}$. From that point onward, it treats the new validator as part of the system and interacts with it according to the protocol.
By round $r^{\prime}+\Delta^{*}$, the new validator will have received all consistency violation reports known to validator $p$ for rounds $\leq r^{\prime}$. For future rounds, the new validator learns of inconsistencies through normal protocol participation. Since this argument holds for any existing validator, it is guaranteed that if an honest validator observes a consistency violation, the joining validator will be aware of the violation.
Recovery Protocols
Our solution ensures that rational validators have no positive utility in misbehaving, thus eliminating the need for a recovery protocol, as consistency violations will not occur in our model. However, in practice, it would be beneficial if the system could recover from failures in the event a rational validator incorrectly estimates its utility. For this, we refer to [4], which –assuming the same threshold of correct validators – describes a protocol that merges conflicting branches, compensates affected users from the stake of misbehaving validators, and reconfigures the validator set accordingly. This is possible as CoBRA guarantees that the stake of the rational validators always exceeds the potential utility (i.e. also economic damage) they can extract. Specifically, in the presence of $i\leq f$ conflicting ledgers, misbehaving validators can double-spend in total up to $(i-1)D$ coins. However, upon equivocation, validators in CoBRA detect the misbehavior of at least $f+1$ validators and slash their stake which amounts to $(f+1)D$.
Assessing Transaction Value
To address this, clients can include an estimated transaction value in their submissions. If the receiver agrees with the estimate, they accept the stake-based finality rule; otherwise, they wait for the full finality threshold $2\Delta^{*}$ (or a strong chain). However, this mechanism introduces a vulnerability: a malicious client can delay the finality of other transactions by regularly issuing, or falsely declaring, high-value transactions. To mitigate this, we introduce a two-tier system in which high-value transactions always finalize at $\Delta^{*}$ (or via strong chains), while only lower-value ones qualify for faster stake-based finality. This makes such attacks costly, as the adversary must submit many high-value transactions to be effective.
Bribing Mechanism
Responsive Synchronous Model
[…]
When applied to partially synchronous protocols, our transformation assumes a model similar to prior work [12, 8]. In addition to the usual partial synchrony assumptions, where message delays are bounded by $\Delta$ after GST, we introduce a larger synchronous bound $\Delta^{*}$ that applies even before GST. This captures the presence of slower, potentially out-of-band communication channels that remain functional during network disruptions or adversarial conditions, with $\Delta^{*}$ possibly orders of magnitude greater than $\Delta$.
8 Related Work
Blockchains usually operate in financial contexts, and thus, their validators which are responsible for their maintenance and security, are typically financially motivated. Consequently, designing incentive-compatible consensus protocols is crucial, particularly given the permissionless nature of blockchains. While production blockchains such as Bitcoin and Ethereum incorporate incentive mechanisms, they are not incentive-compatible [19, 20, 21, 22, 23].
In the PoW setting, the state-of-the-art incentive-compatible blockchain protocol is FruitChains [24], which remains secure as long as rational miners do not form a coalition that exceeds $1/2$ of the total mining power. Several subsequent works have proposed diverse incentives mechanisms for PoW blockchains to optimize specific metrics, e.g., latency [25].
While PoW blockchains continue to evolve, environmental concerns and scalability limitations have driven the blockchain community towards Proof of Stake (PoS) blockchains. Recent research has particularly focused on accountable PoS systems [26, 27, 28, 29, 30], where misbehaving participants can be identified and penalized.
Closely related works that adopt a mixed Byzantine-rational model similar to ours include BAR [31] and Trap [5]. Trap [5], introduces a baiting game to hold misbehaving nodes accountable, solving one-shot consensus in the partial synchronous model for $n>\max\left(\frac{3}{2}k+3f,2(k+f)\right)$. In contrast, our work focuses on multi-shot consensus to realize an SMR protocol. Our solution achieves better resiliency while tolerating up to $f+k<\frac{2}{3}$, compared to Trap’s limit of $f+k<\frac{1}{2}$. While our protocol relies on synchrony, we prove that achieving SMR in a partially synchronous network is impossible under our model. On the other hand, BAR is designed for general-purpose state machine replication and does not address the economic incentive structures or adversarial dynamics specific to payment systems. As such, it is not directly applicable to our work, which explicitly tackles rational deviations in the context of decentralized payment protocols, providing guarantees to validators and clients.
The most closely related works are BLR24 [4] and ZLB [6], both of which investigate economic safety in payment systems. BLR24 examines economic safety in PoS blockchains with the goal to determine when slashing mechanisms can effectively deter misbehaving validators from launching equivocation attacks while ensuring honest validators retain their stake. They prove that slashing is impossible when more than $2f$ out of $13f+$ 1 validators misbehave, and propose a slashing mechanism when at most $2f$ do. Although our bounds align, we also provide resilience up to $2f$ Byzantine and rational nodes, our focus is fundamentally different. BLR24 [4] does not account for liquid coin transfers during SMR execution or explicitly model profit generation beyond slashing deterrents.
ZLB provides probabilistic economic safety in the presence of up to $\frac{5}{9}$ adversarial nodes, where less than $\frac{2}{3}$ can be Byzantine while the rest follow the a-b-c model [32]. Although ZLB supports the partially synchronous model, it does so probabilistically by leveraging a random beacon to construct random committees. In contrast, our approach provides stronger, deterministic guarantees for transaction finality, ensuring that any transaction accepted by clients is never reverted.
Unlike BLR24 [4] and ZLB [6], we introduce a finer classification of adversaries: rational validators, who optimize profit strictly within the PoS framework (e.g., via double-spending), and Byzantine validators, who may have external incentives. Under this model, we prove that finalized transactions will never be reverted, ensuring that rational validators find equivocation unprofitable and that Byzantine adversaries alone cannot cause equivocation. To achieve this stronger finality guarantee, we design novel confirmation rules for both validators and clients, explicitly considering liquid coin transfers and block signature counts within a bounded time window.
Several recent works explore recovery protocols for when the $\frac{1}{3}$ fault threshold is exceeded, as well as the provision of cryptoeconomic safety guarantees for clients [7, 33, 9]. These findings are complementary to ours. Our solution ensures that rational validators have no incentive to misbehave, thereby preventing consistency violations and eliminating the need for recovery protocols within the scope of our model. Nevertheless, in practice, incorporating recovery mechanisms could be beneficial to handle failures caused by unforeseen validator behavior that falls outside the assumptions of our rational model.
Finally, our impossibility results under partial synchrony align with [34, 4], which show that Byzantine validators can cause safety violations and avoid slashing before cashing out once they exceed the security threshold. We extend these findings by showing that, even when the Byzantine threshold is not exceeded, the attack remains feasible and profitable if just one honest validator is replaced by a rational one. Moreover, our finality impossibility (Theorem 3) demonstrates that such attacks remain feasible and profitable for protocols finalizing blocks before a large synchronous bound elapses.
Appendix A Rational attacks
In this section, we illustrate the challenges of extending $(n,f)$–resilient SMR protocols towards $(n,k,f)$–resiliency for $k\geq 1$.
Safety attacks
The attack. Rational validators can collude with adversarial validators to create two conflicting blocks $b_{1}$ and $b_{2}$ with valid inclusion proofs, where they spend $\alpha$ coins to two distinct subsets of clients $C_{1}$ and $C_{2}$. For clients in $C_{1}$ (or $C_{2}$), the $\alpha$ coins have not been spent yet, and therefore, rational validators can spend them again. In that case, we say that rational validators double-spend $\alpha$ coins. We present this attack in Lemma 8, showing that $(n,f)$-resilient blockchains are vulnerable to double-spending.
Lemma 8 (Double-Spending Attack).
Consider a system of $n=f+k+h$ validators, where $f$ are adversarial, $k$ are rational, and $h$ are correct validators. Let $\Delta^{*}$ denote the maximum message delay (either known or not before execution). For any $q\leq f+k+\lfloor h/2\rfloor$ and $f+k>1$ and any q–commitable protocol $\Pi$ that is $(n,f)$-resilient:
  1. 1.
    $\Pi$ is vulnerable to double-spending attacks,
  2. 2.
    if $\Pi$ is broadcast based, any equivocation is witnessed by all validators within $2\Delta^{*}$.

### v3

@@ B.1. Lemmas for proving safety
To complete the withdrawal, a second transaction $R_{p,s}$ must be committed to $L_{i}$ at a time $t^{\prime}+2\Delta^{*}$. Since $L_{i}$ is the first ledger to include $R_{p}$, the same condition applies to any other conflicting ledger, namely the withdrawal can only be finalized after $t^{\prime}+2\Delta^{*}$. However, since correct validators broadcast valid certificates, every correct validator will witness all conflicting ledgers at some time $t^{*}\leq t+\Delta^{*}\leq t^{\prime}+2\Delta^{*}$. As a result, they will detect equivocation and blacklist $p$, preventing the finalization of $R_{p,s}$. Thus, $p$’s withdrawal cannot succeed in any of the conflicting ledgers.
Lemma 0.
Assume that correct validators employ the Strong Chain finalization rule and at least one coalition of validators forks the system. The coalition double spends up to $fD$ coins and any rational validator participating in the attack can double spend at most $D$ coins.
Proof.
Similar to Lemma B.2, we will fix the longest ledger at the time of the equivocation and upper bound the total amount of coins in conflicting blocks that are finalized by at least one validator during $\Delta^{*}$.
Case 1: Every correct validator finalizes up to $D$. The proof is similar to Lemma B.2.
Case 2: At least one correct validator $p_{i}$ monitors an $i$-strong chain $L_{p_{i}}$ with $i>f/4$. Since $1+i$ correct validators monitor $L_{p_{i}}$, there can be at most fif-i additional conflicting ledgers. We assume the worst case, which is that fif-i conflicting ledgers exist. We fix any rational validator $p$ engaging in the attack, participating in fmif-m-i conflicting ledgers $L_{i\in[f-m-i]}$, with $m\in[f-i]$.
For each $i\in[f-m-i]$, at most $m+1$ correct validators and therefore at least fmf-m rational validators monitor ledger $L_{i}$. We prove that by contradiction. Assume there is a ledger $L_{i}$ with $m+2$ correct validators. At least $i+1$ correct validators monitor $L_{p_{i}}$, at least $m+2$ correct validators monitor $L_{i}$ and at least one correct validator exists in the rest fmi1f-m-i-1 conflicting ledgers $L_{j\neq i}$. In total, this accounts for $f+2$ correct validators, leading to a contradiction. Furthermore, since there are at most $m+1$ correct validators per ledger, the total amount of coins finalized by correct validators on any ledger is at most $\frac{f}{f-m}D$.
Finally, there are fmif-m-i conflicting ledgers with at most fmf-m rational validators spending up to $\frac{f}{f-m}D$ per ledger. So, the amount of double-spent coins is bounded by $max_{\{m\in[f-i]\}}\frac{(f-m-i)f}{(f-m)^{2}}D$ which is bounded by $D$ for $i\in\{\frac{f}{4},...,f\}$. We can show that, by taking the continuous extension of the function $F(m)=\frac{(f-m-i)f}{(f-m)^{2}}D$ for $m\in[0,f-i]$ and some constant $i\in\{\lfloor\frac{f}{4}+1\rfloor,...,f\}$, which takes its maximum value for m=f2im=f-2i. Since the function takes maximum value at an integer value, the maximum will be the same for the discrete function and thus $\frac{f}{4i}D\leq D$.
Lemma 0.
At any round, there can only exist a unique strongest chain.
Proof.
For the sake of contradiction, assume that at some round $r$, there exist two conflicting chains $T_{1}$, $T_{2}$ that both form a strongest chain. That means that $T_{1}$ is an $i_{1}$-strong chain for some $i_{1}>\frac{f-1}{2}$ and $T_{2}$ is an $i_{2}$-strong chain for some $i_{2}>\frac{f-1}{2}$. Therefore, there exist two distinct subsets $H_{1}$ and $H_{2}$ that consist of $1+i_{1}$ correct validators participating in $T_{1}$ and $1+i_{2}$ correct validators participating in $T_{2}$ respectively. However, $1+i_{1}+1+i_{2}>f+1$. Since there are $f+1$ only correct validators in the system, we reached a contradiction. ∎
Lemma 0.
Assume that correct validators employ the Strongest Chain finalization rule and at least one coalition of validators forks the system. The coalition double spends up to $fD$ coins and any rational validator participating in the attack can double-spend at most $D$ coins.
Proof.
We will prove that when validators employ the strongest chain finalization rule, rational validators participating in coalitions can double-spend up to $D$ coins. Similar to Lemma B.2, we fix the longest ledger at the time of the equivocation and provide an upper bound on the total amount of coins included in conflicting blocks that are finalized by at least one validator within the interval $\Delta^{*}$.
Case 1: There is no correct validator $p_{i}$ monitoring an $i_{1}$-strong chain with $i_{1}>(f+1)/2$. If every correct validator finalizes up to $D$, the proof follows from Lemma B.2. If at least one correct validator monitors an $i_{1}$-strong chain with $i_{1}>f/4$, the proof follows from Lemma B.4.
Case 2: At least one correct validator $p_{i}$ monitors an $i_{1}$-strong chain with $i_{1}>(f+1)/2$. According to Lemma B.5, only one strongest chain can exist. Thus we have the following cases.
Case 2a: Every correct validator in a conflicting ledger, finalizes up to $D$. A number of $1+i_{1}$ correct validators monitor $L_{p_{i}}$, and thus there are up to $f-i_{1}$ extra conflicting ledgers. Assume that there are $f-i_{1}$ extra conflicting ledgers and let us any rational validator participating in $f-m-i_{1}$ of the conflicting ledgers for $m\in[f-i]$. In each fork, there are at most $m+1$ correct validators and at least fmf-m rational validators (with the same analysis done in Lemma B.2). Therefore, rational misbehaving validators can double-spend up to $max_{\{m\in[f-i_{1}]\}}\frac{(f-m-i_{1})}{(f-m)}D\leq{m\in[f-i_{1}]\}}\frac{(f-m)}{(f-m)}D=D$.
Now, we fix any rational validator $p$ participating in $L_{p_{i}}$ and in $i_{3}-m$ of the conflicting ledgers $L_{i\in[i_{3}-m]}$ for $m\in[i_{3}]$. We will upper bound the coins that $p$ can spend in these conflicting blocks. First, in $L_{p_{j}}$, there are finalized up to $\frac{f}{(f-m)(f-i_{2})}D<A(m)=\frac{f}{(f-m)(f-\frac{f-1}{2})}D$. For the conflicting ledgers $L_{i\in[i_{3}-m]}$, we can show that there are at most $m+1$ correct validators and at least fmf-m rational validators per fork, with the same analysis done in B.2. Thus, each validator can double-spend up to $\frac{f+1-i_{1}^{*}-i_{2}^{*}-m}{f-m}D=(1+\frac{1-i_{1}^{*}-i_{2}^{*}}{f-m})D\leq B(m)=1-\frac{3(f+2)}{4(f-m)})D$. Now we need to upper bound $C(m)=A(m)+B(m)=(1-\frac{3f^{2}+f+6}{4(f-m)(f+1)})D<D$ since $m\in[f+1-i_{1}^{*}-i_{2}^{*}]$ and $f>0$.
@@
Finally, in Lemmas B.7B.8, we conclude that rational validators have no incentive to engage in forking attacks, as any potential gain does not exceed their staked amount, which is subject to slashing.
@@
Let us fix any rational validator $p$ and compute its utility to join coalitions forking the system. In Lemmas B.2B.4B.6, we showed that by performing the attack, $p$ can double-spend at most $D$ coins when validators employ i) finalization rule presented in 4, ii) strong ledger finalization rule, iii) strongest finalization rule respectively. According to Lemma B.3, $p_{j}$ cannot withdraw its stake $D$ before the conflicting ledgers are detected. Moreover, upon detecting equivocation the system halts and thus misbehaving validators cannot utilize further bribes. Therefore, for any strategy $s_{p}\in S^{p}_{\mathsf{f}}$ and $s_{-p}\in S^{-p}_{\mathsf{f}}$, the utility of $p$ is $u_{p}(s_{p},s_{-p})\leq 0$, while following the protocol yields non-negative utility.
To show that following the protocol weakly dominates the strategy of forking the system, we will construct a configuration where $p$ chooses a strategy $s_{p}\in S^{p}_{\mathsf{f}}$ and fix the rest strategies $s_{-p}\in S^{-p}_{\mathsf{f}}$ such that $u_{p}(s_{p},s_{-p})<0$ (this is sufficient since following the protocol has non-negative utility). To this end, suppose that $p$ participates in a coalition $\mathcal{F}$ consisting of $f$ Byzantine and $\lceil\frac{f+1}{2}\rceil$ rational validators. Moreover, consider a subset $H_{1}$ of $\lceil\frac{f+1}{2}\rceil$ correct validators, and a subset $H_{2}$ distinct to $H_{1}$ that consists of $\lfloor\frac{f+1}{2}\rfloor$ correct validators. Validators in $\mathcal{F}$ construct to conflicting ledgers $T_{1}$ and $T_{2}$ as described in Lemma A.1. The utility of $p$ in that case is $u_{p}(s_{p},s_{-p})\leq\frac{D}{\lceil\frac{f+1}{2}\rceil}-D<0$.
@@
For the sake of contradiction assume that two distinct correct validators $p_{1}$, $p_{2}$ finalize the conflicting blocks $b_{1}$ and $b_{2}$ respectively. Validator $p_{1}$ finalized $b_{1}$ after witnessing the respective valid certificate $QC(b_{1})$, and $p_{2}$ finalized $b_{2}$ after witnessing $QC(b_{2})$. By quorum intersection there exists a validator $p$ that is not Byzantine and has voted in both $QC(b_{1})$ and $QC(b_{2})$. Since correct validators do not equivocate, $p$ must be rational. However, according to Lemma B.7, for rational validators, following the protocol weakly dominates participating in a coalition that forks the system. So, $p$ will not sign for conflicting blocks and thus we reached a contradiction.
B.2. Lemmas for proving liveness
Since $\Pi$ is $(n,f)$-resilient and $f+1+k\geq n-f$, liveness is ensured as long as rational validators participate in liveness-preserving actions during some intervals. In Lemma B.9, we show that there exist windows during which it is a dominant strategy for rational validators to follow the protocol specifications. Building on this, Lemma B.10 concludes that some correct validator will successfully propose a block that gathers a valid certificate and finality votes.
@@ B.2. Lemmas for proving liveness
Forking. From Lemma B.7, any strategy in the strategy profile $S^{p}_{f}$ (i.e., where $p$ actively contributes to a fork), is weakly dominated by following the protocol. Therefore, no rational validator will engage in forking attacks.
@@
Suppose the adversary is able to bribe at least one rational validator in every window $W_{m}$ following round $r$. In such windows, at most $2f$ validators can successfully propose valid blocks. Since participation rewards are distributed only when at least 2f+12f{+}1 distinct validators succeed, no rewards are paid out in these bribed windows. Because the adversary’s budget is finite, it is eventually depleted and the adversary cannot bribe any rational validator anymore. Therefore, there exists a future window $W_{k}$ (starting at some round $r^{\prime}\geq r$) such that, for any validator $p$, any strategy $s_{p}\in S_{p}$, and any $s_{-p}\in S_{-p}$ under which $p$ abstains from liveness actions during $W_{k}$, we have $u_{p}(s_{p},s_{-p})\leq 0$.
To show that following the protocol during $W_{k}$ is a weakly dominant strategy, fix any rational validator $p$, and consider two strategies: $s_{1}$, where $p$ follows the protocol during $W_{k}$, and $s_{2}$, where $p$ abstains from liveness actions during $W_{k}$ (we have already established the result for forking in B.7). Let $s_{-p}$ be a strategy profile in which at least $2f$ other validators follow the protocol during $W_{k}$. In the combined strategy profile $s_{a}=(s_{1},s_{-p})$, at least $2f+1$ validators follow the protocol, satisfying the reward condition, and $p$ receives the corresponding participation rewards. In contrast, under $s_{b}=(s_{2},s_{-p})$, only $2f$ validators follow the protocol, so no rewards are distributed. Hence, $u^{W_{k}}_{p}(s_{a})>u^{W_{k}}_{p}(s_{b})$.
@@
  2. (2)
    $2f+1$ finality votes for $b$ or a subsequent block are included in blocks with valid certificates.
@@
According to Lemma B.9, for any round $r$, there exists a round $r^{\prime}\geq r$ such that a time window $W_{k}$ begins, during which it is a weakly dominant strategy for rational validators to follow the protocol. Consequently, some block $b$ proposed by an honest validator during $W_{k}$ will obtain a valid quorum certificate $QC(b)$ and receive at least $2f+1$ finality votes (either for $b$ or for one of its descendants).
B.3. Final Theorems
Theorem B.11.
Validators running the composition of an $(3f+1,f)$-resilient, (2f+1)(2f+1)–commitable SMR protocol $\Pi$ and CoBRA agree on a total order in all executions with $f^{*}\leq f$ and $k\leq 2f-f^{*}$.
@@ B.3. Final Theorems
Safety. In Lemma B.8, we show that validators will never finalize conflicting blocks.
Liveness. Consider any transaction $tx$ witnessed by all correct validators at some round $r$. According to Lemma B.10, there is a round $r^{*}\geq r$ where a correct validator $p$ will successfully propose a valid block $b$. If $tx$ has not already been included in a valid block proposed before round $r^{*}$, then, $p$ will include $tx$ at block $b$.
Theorem B.12.
The composition of an $(3f+1,f)$-resilient, (2f+1)(2f+1)–commitable SMR protocol $\Pi$ and CoBRA satisfies certifiability in all executions with $f^{*}\leq f$ and $k\leq 2f-f^{*}$.
@@
To this end, consider the inclusion proof $\pi_{tx}$ for the transaction $tx$. Client $c$ has accepted $\pi_{tx}$ only after receiving a proof showing the inclusion of $tx$ in a block $b$, where $b$ has $2f+1$ finality votes. At least one correct validator $p$ has committed a finality vote for $b$. Validator $p$ commits finality votes only for blocks included in its transaction ledger.
Client-liveness. For every transaction $tx$ committed to a block $b$ accepted by all correct validators, any correct validator $p_{i}$ can construct a proof $\pi$ as follows. Validator $p_{i}$ sends an inclusion proof $\pi_{tx}$ of $tx$ in $b$ to a client $c$, along with a proof $\pi_{f}(b)$ consisting of $2f+1$ finality votes for $b$ or subsequent block, which is feasible according to Lemma B.10. Any client $c$ will accept the proof $\pi$ by definition of the client confirmation rule.
Theorem B.13.
The composition of CoBRA and a (2f+1)(2f\!+\!1)–committable protocol $\Pi$ is an $(n,k,f^{*})$-resilient SMR protocol in all executions with $f^{*}\leq f$ and $k\leq 2f-f^{*}$.
@@
Validators running the composition of CoBRA and $\Pi$ agree in a total order according to Theorem B.11. Moreover, in Theorem B.12 we show that the composition of CoBRA and $\Pi$ satisfies certifiability.
Theorem B.14.
The composition of an $(3f+1,f)$-resilient, (2f+1)(2f+1)–commitable SMR protocol $\Pi$ and CoBRA regains economic restitution with recovery parameter $\Delta_{R}=4\Delta^{*}$ in all executions with $f^{*}<5n/9$ and, $k<2n/3-f^{*}$.